OpenStudy (anonymous):

$42 over 2 years at an interest rate of 3%

OpenStudy (anonymous):

• P is the principal amount, $42.00. • r is the interest rate, 3% per year, or in decimal form, 3/100=0.03. • t is the time involved, 2....year(s) time periods. • So, t is 2....year time periods. To find the simple interest, we multiply 42 × 0.03 × 2 to get that: The interest is: $2.52 Usually now, the interest is added onto the principal to figure some new amount after 2 year(s), or 42.00 + 2.52 = 44.52. For example: • If you borrowed the $42.00, you would now owe $44.52 • If you loaned someone $42.00, you would now be due $44.52 • If owned something, like a $42.00 bond, it would be worth $44.52 now.

OpenStudy (anonymous):

It's either 400 500 600 or 700

OpenStudy (wolf1728):

The formula is: present value = principal * (1 + rate) ^ years present value = principal * (1.03) ^ 2 present value = 42 * 1.0609 present value = 44.5578 present value = $44.56 (A little higher than the previous answer.)

OpenStudy (anonymous):

So what would be the answer 400 590 600 or 700 @wolf1738

OpenStudy (wolf1728):

$44.56

OpenStudy (anonymous):

either mine or wolf1728 is correct it all depends on how u rounded your decimals

OpenStudy (anonymous):

The answers are either 400 500 600 or 700 it's muni tole choice

OpenStudy (anonymous):

Multiple choice

OpenStudy (anonymous):

@wolf1728

OpenStudy (anonymous):

its 700

OpenStudy (wolf1728):

Where does 700 come into the picture?

OpenStudy (anonymous):

Match the information about interest earned with the principal amount invested . Remember the equation for calculating simple interest is I=prt.

OpenStudy (wolf1728):

princ = 42 int = .03 years = 2 total = 44.52 and something in that computes to 700?

OpenStudy (anonymous):

So what would be the answer wolf?

OpenStudy (wolf1728):

I don't know. For one thing I didn't know it was a multiple choice question. The question was $42 over 2 years at an interest rate of 3% and I just assumed it was to calculates the value after 2 years.